OpenStudy (anonymous):

On the blueprint of a house, 48 millimeters represents 9 meters. The length of the living room is 32 millimeters on the blueprint. What is the actual length of the living room?

OpenStudy (anonymous):

6 m

OpenStudy (anonymous):

the fraction of 48 over 32 has to be equal to the fraction 9 over x where x represents the length of the living room

OpenStudy (anonymous):

\[\frac{ 48 }{ 9}=\frac{ 32}{ x }\] 48x=32(9) solve for x

OpenStudy (anonymous):

both ways get the same asnwer